# PSMGD: Periodic Stochastic Multi-Gradient Descent for Fast Multi-Objective Optimization

### v1

@@ Abstract
Multi-objective optimization (MOO) lies at the core of many machine learning (ML) applications that involve multiple, potentially conflicting objectives (e.g., multi-task learning, multi-objective reinforcement learning, among many others). Despite the long history of MOO, recent years have witnessed a surge in interest within the ML community in the development of gradient manipulation algorithms for MOO, thanks to the availability of gradient information in many ML problems. However, existing gradient manipulation methods for MOO often suffer from long training times, primarily due to the need for computing dynamic weights by solving an additional optimization problem to determine a common descent direction that can decrease all objectives simultaneously. To address this challenge, we propose a new and efficient algorithm called Periodic Stochastic Multi-Gradient Descent (PSMGD) to accelerate MOO. PSMGD is motivated by the key observation that dynamic weights across objectives exhibit small changes under minor updates over short intervals during the optimization process. Consequently, our PSMGD algorithm is designed to periodically compute these dynamic weights and utilizes them repeatedly, thereby effectively reducing the computational overload. Theoretically, we prove that PSMGD can achieve state-of-the-art convergence rates for strongly-convex, general convex, and non-convex functions. Additionally, we introduce a new computational complexity measure, termed backpropagation complexity, and demonstrate that PSMGD could achieve an objective-independent backpropagation complexity. Through extensive experiments, we verify that PSMGD can provide comparable or superior performance to state-of-the-art MOO algorithms while significantly reducing training time.
@@
Extended versionhttps://aaai.org/example
@@ 1 Introduction
Just as the human world is full of diverse and potentially conflicting goals, many machine learning paradigms are also multi-objective, such as multi-task learning (Caruana 1997; Sener and Koltun 2018), multi-objective reinforcement learning (Hayes et al. 2022), learning-to-rank (Mahapatra et al. 2023; Mahapatra, Dong, and Momma 2023), among others. At the core of these machine learning paradigms lies multi-objective optimization (MOO), which aims to optimize multiple potentially conflicting objectives simultaneously. Formally, MOO can be mathematically cast as:
@@
The most common approach for handling MOO is to optimize a weighted average of the multiple objectives, also known as static/unitary weighting or linear scalarization (Kurin et al. 2022; Xin et al. 2022). This approach transforms MOO into a single-objective problem, for which there are various off-the-shelf methods available. The linear scalarization approach has a fast per-iteration runtime but suffers from slow convergence and poor performance due to potential conflicts among objectives. To address this problem, there has been a surge of interest in recent years in developing gradient manipulation algorithms for MOO (Kendall, Gal, and Cipolla 2018; Chen et al. 2018; Yu et al. 2020; Chen et al. 2020; Javaloy and Valera 2021; Liu et al. 2021a; Liu and Vicente 2021; Chen et al. 2024; Xiao, Ban, and Ji 2024). The key idea is to calculate dynamic weights to avoid conflicts and find a direction $\mathbf{d}({\mathbf{x}})$ that decreases all objective functions simultaneously. This common descent vector $\mathbf{d}({\mathbf{x}})$ is often determined by solving an additional optimization problem that involves all objective gradients. Updating the model with the common descent direction $\mathbf{d}$ ensures a decrease in every objective. While these approaches show improved performance, the extra optimization process requires a substantial amount of time. This is particularly evident when the number of objectives is large and the model parameters are high-dimensional, which could significantly prolong the MOO training process. Existing empirical evaluations consistently demonstrate that gradient manipulation algorithms typically necessitate much longer training times (Kurin et al. 2022; Liu et al. 2024) (also shown in Sec 4). To this end, a natural question arises: Can we design efficient gradient manipulation algorithms for fast MOO with theoretical guarantees?
In this paper, we present Periodic Stochastic Multi-Gradient Descent (PSMGD), a simple yet effective algorithm to accelerate the MOO process. Our PSMGD is motivated by the key observation that dynamic weights across objectives exhibit small changes under minor updates over short intervals. Thus, our PSMGD algorithm is designed to periodically compute these dynamic weights and utilizes them repeatedly, thereby effectively reducing the computational overload and achieving a faster MOO training process. Our contributions are summarized as follows:
@@
  • Built upon the key observation, we introduce PSMGD, a new and efficient gradient manipulation method algorithm. By calculating dynamic weights across objectives infrequently, PSMGD significantly alleviates computational burden and reduces training time.
  • We conduct a theoretical analysis of PSMGD for strongly convex, general convex, and non-convex functions, demonstrating that PSMGD achieves state-of-the-art convergence rates comparable to existing MOO methods (Table 1). We also introduce a new computational complexity measure, backpropagation (BP) complexity, to quantify computational workload, and further show that PSMGD can achieve an objective-independent backpropagation complexity. Compared with existing MOO methods, this shows a superior BP complexity with the linear speedup with respect to the number of objectives.
@@ 2 Related works
The MOO problem, as stated in Eq. 1, has a long and rich history that dates back to the 1950s. MOO algorithms can be broadly categorized into two main groups. The first category comprises gradient-free methods, such as evolutionary MOO algorithms and Bayesian MOO algorithms (Zhang and Li 2007; Deb et al. 2002; Belakaria et al. 2020; Laumanns and Ocenasek 2002). These methods are more suitable for small-scale problems but are less practical for high-dimensional MOO models, such as deep neural networks. The second category is the gradient-based approach by utilizing (stochastic) gradients (Fliege and Svaiter 2000; Désidéri 2012; Fliege, Vaz, and Vicente 2019; Liu and Vicente 2021), making them more practical for high-dimensional MOO problems. In this work, we primarily focus on gradient-based approaches to solve MOO in high-dimensional deep-learning models. In what follows, we provide a brief overview of two typical approaches for gradient-based methods for MOO: linear scalarization (LS) and gradient manipulation methods.
@@
1) LS: A straightforward approach for MOO 1 is to transform it into a single objective problem by using pre-defined weights $\{\lambda_{s},s\in[s]\}$: $\min_{{\mathbf{x}}\in\mathcal{D}}\sum_{s\in[S]}\lambda_{s}f_{s}({\mathbf{x}}).$ Thanks to the LS, one can leverage many existing and familiar single-objective methods, such as stochastic gradient descent, for training at each iteration. However, in LS, it is not uncommon to see conflicts among multiple objectives during the optimization process, i.e., $\nabla f_{s}({\mathbf{x}}),\nabla f_{s^{\prime}}({\mathbf{x}})\right><0$. This implies that the update using static weights may decrease some objectives, while inevitably increasing others at the same time and leading to slow convergence and poor performance.
@@
2) Gradient Manipulation Methods: A popular alternative is to dynamically weight gradients across objectives to avoid such conflicts and obtain a direction to decrease all objectives simultaneously. For example, multiple gradient descent algorithm (MGDA) (Fliege and Svaiter 2000) is seeking to find the common descent direction by solving the following optimization problem given each objective’s gradient: $(\mathbf{d},\beta)\in\operatorname*{arg\,min}_{\mathbf{d}\in\mathbb{R}^{S},\beta\in\mathbb{R}}\beta+\frac{1}{2}\|\mathbf{d}\|^{2},s.t.,\nabla f_{s}({\mathbf{x}})^{T}\mathbf{d}-\beta\leq 0,\forall s\in[S].$ By doing so, if ${\mathbf{x}}$ represents a first-order Pareto stationary point, then $(\mathbf{d},\beta)=(\mathbf{0},0)$. Otherwise, we can find the $\mathbf{d}$ such that $\nabla f_{s}({\mathbf{x}})^{T}\mathbf{d}\leq\beta<0,s\in[S]$ as the solution. Using such a non-conflicting direction $\mathbf{d}$ to update the model, i.e., ${\mathbf{x}}\leftarrow{\mathbf{x}}+\eta{\mathrm{d}}$ where $\eta$ is the learning rate, has been shown to achieve better performance in practice (Sener and Koltun 2018; Liu et al. 2021a). Following this token, many gradient manipulation methods attempt to obtain a common descent direction through various formulations and solutions. For example, GradDrop (Chen et al. 2020) randomly dropped out highly conflicted gradients, RotoGrad (Javaloy and Valera 2021) rotated objective gradients to alleviate the conflict, and many other methods exploring similar principles (Kendall, Gal, and Cipolla 2018; Chen et al. 2018; Yu et al. 2020; Liu et al. 2021a; Liu and Vicente 2021; Chen et al. 2024; Xiao, Ban, and Ji 2024). However, due to the additional optimization process required to generate such $\mathbf{d}$, gradient manipulation methods often incur an expensive per-iteration cost, thereby prolonging training time. In this work, we propose a new algorithm, PSMGD, which combines the advantages of both approaches and offers a fast per-iteration runtime, a fast overall convergence speed, and improved model performance.
@@ 3.2 A pedagogical example
Observation 2: The dynamic weights calculated by MGDA converge quickly and exhibit stability. As shown in Figure 1(d), we visualize the dynamic weights for the Fonseca problem. In this simple problem, the weights converge within 5 iterations, after which they remain unchanged. In more complex problems, such as multi-task learning using deep learning (see Sec B in Appendix), the dynamic weights calculated by gradient manipulation methods also demonstrate stability. We verified stability by examining two scales, epoch and iteration. The weights show consistent behavior with a stable average value and low variability across both scales, epoch and iteration, as depicted in Figure 6(a) and 6(b). This observation suggests a new approach to develop more efficient MOO algorithms. Specifically, we can calculate and update the dynamic weights periodically in a more lazy and thus efficient manner.
@@ 3.4 Convergence analysis
Before providing the convergence analysis, we have the following assumptions:
@@
There exists a constant $L>0$ such that $\|\nabla f_{s}({\mathbf{x}})-\nabla f_{s}({\mathbf{y}})\|\leq L\|{\mathbf{x}}-{\mathbf{y}}\|,\forall{\mathbf{x}},{\mathbf{y}}\in\mathbb{R}^{d},s\in[S]$.
Assumption 3.2 (Unbiased stochastic gradient with bounded variance).
We assume the stochastic gradient estimation of each objective function is unbiased with bounded variance.

### v2

@@ Abstract
Multi-objective optimization (MOO) lies at the core of many machine learning (ML) applications that involve multiple, potentially conflicting objectives. Despite the long history of MOO, recent years have witnessed a surge in interest within the ML community in the development of gradient manipulation algorithms for MOO, thanks to the availability of gradient information in many ML problems. However, existing gradient manipulation methods for MOO often suffer from long training times, primarily due to the need for computing dynamic weights by solving an additional optimization problem to determine a common descent direction that can decrease all objectives simultaneously. To address this challenge, we propose a new and efficient algorithm called Periodic Stochastic Multi-Gradient Descent (PSMGD) to accelerate MOO. PSMGD is motivated by the key observation that dynamic weights across objectives exhibit small changes under minor updates over short intervals during the optimization process. Consequently, our PSMGD algorithm is designed to periodically compute these dynamic weights and utilizes them repeatedly, thereby effectively reducing the computational overload. Theoretically, we prove that PSMGD can achieve state-of-the-art convergence rates for strongly-convex, general convex, and non-convex functions. Additionally, we introduce a new computational complexity measure, termed backpropagation complexity, and demonstrate that PSMGD could achieve an objective-independent backpropagation complexity. Through extensive experiments, we verify that PSMGD can provide comparable or superior performance to state-of-the-art MOO algorithms while significantly reducing training time.
@@
Extended versionhttps://arxiv.org/abs/2412.10961
@@ 1 Introduction
Just as the human world is full of diverse and potentially conflicting goals, many machine learning paradigms are also multi-objective, such as multi-objective reinforcement learning (Hayes et al. 2022), multi-task learning (Caruana 1997; Sener and Koltun 2018), and learning-to-rank (Mahapatra et al. 2023). At the core of these machine learning paradigms lies multi-objective optimization (MOO), which aims to optimize multiple potentially conflicting objectives simultaneously. Formally, MOO can be mathematically cast as:
@@
The most common approach for handling MOO is to optimize a weighted average of the multiple objectives, also known as linear scalarization (Kurin et al. 2022; Xin et al. 2022). This approach transforms MOO into a single-objective problem, for which there are various off-the-shelf methods available. The linear scalarization approach has a fast per-iteration runtime but suffers from slow convergence and poor performance due to potential conflicts among objectives. To address this problem, there has been a surge of interest in recent years in developing gradient manipulation algorithms (Kendall, Gal, and Cipolla 2018; Chen et al. 2018; Yu et al. 2020; Chen et al. 2020; Javaloy and Valera 2021; Liu and Vicente 2021; Chen et al. 2024; Xiao, Ban, and Ji 2024). The key idea is to calculate dynamic weights to avoid conflicts and find a direction $\mathbf{d}({\mathbf{x}})$ that decreases all objectives simultaneously. This common descent vector $\mathbf{d}({\mathbf{x}})$ is often determined by solving an additional optimization problem that involves all objective gradients. Updating the model with the common descent direction $\mathbf{d}$ ensures a decrease in every objective. While these approaches show improved performance, the extra optimization process requires a substantial amount of time. This is particularly evident when the number of objectives is large and the model parameters are high-dimensional, which could significantly prolong the MOO training process. Existing empirical evaluations consistently demonstrate that gradient manipulation algorithms typically necessitate much longer training times (Kurin et al. 2022; Liu et al. 2024) (also shown in Sec 4). To this end, a natural question arises: Can we design efficient gradient manipulation algorithms for fast MOO with theoretical guarantees?
In this paper, we present Periodic Stochastic Multi-Gradient Descent (PSMGD), a simple yet effective algorithm to accelerate the MOO. Our PSMGD is motivated by the key observation that dynamic weights across objectives exhibit small changes under minor updates over short intervals. Thus, it is designed to periodically compute these dynamic weights and utilizes them repeatedly, thereby effectively reducing the computational overload and achieving a faster MOO training process. Our contributions are summarized as follows:
@@
  • Built upon the key observation, we introduce PSMGD, a new and efficient gradient manipulation method. By calculating dynamic weights infrequently, it significantly alleviates computational burden and reduces training time.
  • We conduct a theoretical analysis of PSMGD for strongly convex, convex, and non-convex functions, demonstrating that PSMGD achieves state-of-the-art convergence rates comparable to existing MOO methods (Table 1). We also introduce a new computational complexity measure, backpropagation (BP) complexity, to quantify computational workload, and further show that PSMGD can achieve an objective-independent backpropagation complexity.
@@ 2 Related works
The MOO problem, as stated in Eq. 1, has a long history that dates back to the 1950s. MOO algorithms can be broadly categorized into two main groups. The first category comprises gradient-free methods, such as evolutionary MOO algorithms and Bayesian MOO algorithms (Zhang and Li 2007; Deb et al. 2002; Belakaria et al. 2020; Laumanns and Ocenasek 2002). These methods are more suitable for small-scale problems but are less practical for high-dimensional models, such as deep neural networks. The second category is the gradient-based approach by utilizing (stochastic) gradients (Fliege and Svaiter 2000; Désidéri 2012; Fliege, Vaz, and Vicente 2019; Liu and Vicente 2021), making them more practical for high-dimensional MOO problems. In this work, we primarily focus on gradient-based approaches to solve MOO in high-dimensional deep-learning models. We first provide a brief overview of two typical approaches: linear scalarization (LS) and gradient manipulation methods.
@@
1) LS: A straightforward approach is to transform it into a single objective problem by using pre-defined weights $\bm{\lambda}$: $\min_{{\mathbf{x}}\in\mathcal{D}}\bm{\lambda}^{\top}\mathbf{F}({\mathbf{x}}).$ Thanks to the LS, one can leverage many existing single-objective methods (e.g., gradient descent), for training at each iteration. However, in LS, it is not uncommon to see conflicts among multiple objectives during the optimization process, i.e., $\nabla f_{s}({\mathbf{x}}),\nabla f_{s^{\prime}}({\mathbf{x}})\right><0$. This implies that the update using static weights may decrease some objectives, while inevitably increasing others at the same time and leading to slow convergence and poor performance.
@@
2) Gradient Manipulation Methods: A popular alternative is to dynamically weight gradients across objectives to avoid such conflicts and obtain a direction $\mathbf{d}$ to decrease all objectives simultaneously. For example, multiple gradient descent algorithm (MGDA) (Fliege and Svaiter 2000) seeks to find the $\mathbf{d}$ by solving the following optimization problem given gradients: $(\mathbf{d},\beta)\in\operatorname*{arg\,min}_{\mathbf{d}\in\mathbb{R}^{S},\beta\in\mathbb{R}}\beta+\frac{1}{2}\|\mathbf{d}\|^{2},s.t.,\nabla f_{s}({\mathbf{x}})^{T}\mathbf{d}-\beta\leq 0,\forall s\in[S].$ By doing so, if ${\mathbf{x}}$ represents a first-order Pareto stationary point, then $(\mathbf{d},\beta)=(\mathbf{0},0)$. Otherwise, we can find the $\mathbf{d}$ such that $\nabla f_{s}({\mathbf{x}})^{T}\mathbf{d}\leq\beta<0,s\in[S]$ as the solution. Using such a non-conflicting direction $\mathbf{d}$ to update the model, i.e., ${\mathbf{x}}\leftarrow{\mathbf{x}}+\eta{\mathrm{d}}$ where $\eta$ is the learning rate, has been shown to achieve better performance in practice (Sener and Koltun 2018; Liu et al. 2021a). Following this token, many gradient manipulation methods attempt to obtain a common descent direction through various formulations and solutions. For example, GradDrop (Chen et al. 2020) randomly dropped out highly conflicted gradients, RotoGrad (Javaloy and Valera 2021) rotated objective gradients to alleviate the conflict, and many other methods exploring similar principles (Kendall, Gal, and Cipolla 2018; Chen et al. 2018; Yu et al. 2020; Liu et al. 2021a; Liu and Vicente 2021; Chen et al. 2024; Xiao, Ban, and Ji 2024). However, due to the additional optimization process required to generate such $\mathbf{d}$, gradient manipulation methods often incur an expensive per-iteration cost, thereby prolonging training time. In this work, we propose a new algorithm, PSMGD, which achieves fast per-iteration and overall convergence, along with enhanced performance.
@@ 3.2 A pedagogical example
Observation 2: The dynamic weights calculated by MGDA converge quickly and exhibit stability. As shown in Figure 1(d), we visualize the dynamic weights for the Fonseca problem. In this simple problem, the weights converge within 5 iterations, after which they remain unchanged. In more complex problems, such as multi-task learning using deep learning (see Sec B in Appendix), the dynamic weights calculated by gradient manipulation methods also demonstrate stability. We verified stability by examining two scales, epoch and iteration. The weights show consistent behavior with a stable average value and low variability across both scales, epoch and iteration, as depicted in Figure 6(a) and 6(b) in Appendix. This observation suggests a new approach to develop more efficient MOO algorithms. Specifically, we can calculate and update the dynamic weights periodically in a more lazy and thus efficient manner.
@@ 3.4 Convergence analysis
$\|\nabla f_{s}({\mathbf{x}})-\nabla f_{s}({\mathbf{y}})\|\leq L\|{\mathbf{x}}-{\mathbf{y}}\|,\forall{\mathbf{x}},{\mathbf{y}}\in\mathbb{R}^{d},s\in[S]$.
Assumption 3.2 (Bounded variance).
We assume the stochastic gradient estimation is unbiased with bounded variance.
@@ 5 Conclusion
Acknowledgments
JL acknowledges the funding from NSF grants CAREER CNS-2110259, CNS-2112471, IIS-2324052, DARPA YFA D24AP00265, ONR grant N00014-24-1-2729, and AFRL grant PGSC-SC-111374-19s. HY acknowledges the funding support from AI Seed Funding and GWBC Award at RIT.